sunisround:

How do I write a polynomial function with a degree of 4 and real coefficients in standard form that has 3, -1, and 1+5i as its zeros?

Falconmaster:

@Vocaloid

sunisround:

Can you help me with this problem?

Vocaloid:

if 1 + 5i is a zero then 1 - 5i must also be a zero [forgot what the rule is called]

sunisround:

okay...

Vocaloid:

so for every zero you are given, write an expression for x that gives you that zero for example, for "3" let (x-3) be part of your final function (since x - 3 = 0 gives you 3 as a zero)

sunisround:

how do i write and expression for x that gives me the complex numbers?

Falconmaster:

Oh Welcome to Questioncove

Vocaloid:

then you have (x - 3) (x + 1) (x - [1+5i]) (x + [1-5i]) multiply these together and you should have a polynomial with degree 4

sunisround:

okay.

sunisround:

thank you.

sunisround:

wait

sunisround:

(x-[1+5i]) gives me what result?

Vocaloid:

made a typo on that last one (x - 3) (x + 1) (x - [1+5i]) (x - [1-5i])

Vocaloid:

anyway, if you multiply these 4 things together you should get the answer

Vocaloid:

(x-[1+5i]) means that x = 1 + 5i is a zero [as stated by the original problem]

sunisround:

how did you get the last two?

sunisround:

ok

Vocaloid:

if x = 1 + 5i = 0 then x - (1+5i) = 0 and x - (1 - 5i) = 0 using that same logic

sunisround:

so (x-[1-5i]) means that...?

Vocaloid:

1 - 5i is one of the zeros, with me so far?

sunisround:

yes, so (x-[1-5i]) is x=1-5i?

Vocaloid:

yes

sunisround:

ok let me solve it then

Vocaloid:

you don't need to solve it

Vocaloid:

you just need to write down all of your zero expressions and then multiply them together

sunisround:

that what i meant (misunderstanding)

sunisround:

but wait why when you typed (x-[1+5i]) and (x-[1-5i]) you used x- and not x+?

Vocaloid:

if 3 is our zero, then we have x - 3 = 0 as our zero expression, right? same logic applies for [1-5i]

Vocaloid:

x - the zero = 0

sunisround:

ok good

sunisround:

can you please help me how do i multiply (x-[1+5i]) and (x-[1-5i])?

Vocaloid:

use foil

sunisround:

ok

sunisround:

how do i multiply -[1+5i] and -[1-5i]?

Vocaloid:

(A-B)(C-D) = AC - AD - BC + BD treat [1+5i] and [1-5i] as your B and D values

sunisround:

ok let me try it will probably last about two minutes for my brain to process please be patient with me.

Vocaloid:

it's perfectly fine, take your time and be careful with your signs ^^

sunisround:

(x^2-2x-3)(x^2-x[1-5i]-x[1+5i] is this correct so far?

sunisround:

i still need the last part the L from FOIL

Vocaloid:

yeah that's what I was going to say^^

sunisround:

but i am confused on that one

Vocaloid:

the last term is just [1 + 5i] * [1-5i] which can be solved using foil or difference of squares

Vocaloid:

(A+B)(A-B) = A^2 - B^2

sunisround:

but what about the negatives that where in front of them?

sunisround:

they are positive

Vocaloid:

the negative signs cancel out

sunisround:

so if it would of had a negative and a positive then it would be a negative right?

Vocaloid:

sure

sunisround:

when dealing with complex numbers?

Vocaloid:

(-[1 + 5i]) * (-[1-5i]) = (-1)(-1)[1 + 5i] * [1-5i] so the signs cancel out

Vocaloid:

the fact that they're complex numbers have nothing to do with this, this is just basic foil rules

Vocaloid:

we're not using the signs inside the complex terms yet

sunisround:

so the result of multiplying them is -9?

Vocaloid:

that's not what I get ^^ check your algebra again

Vocaloid:

[1 + 5i] * [1-5i] = 1^2 - (5i)^2 = ?

sunisround:

ok let me mind process it about 2 mins

Vocaloid:

remember that i^2 = -1

sunisround:

what does -(5i)^2 represent?

Vocaloid:

the quantity 5i squared, then multiplied by -1

sunisround:

so that means you got -5i and -5i for the OI part of the FOIL method?

Vocaloid:

one of those should be + 5i

sunisround:

because i got -5i and + 5i

Vocaloid:

good, those cancel out so we only have the first and last terms

sunisround:

I got 1-5i+5i-10i^2

Vocaloid:

check your last term again

Vocaloid:

5i * 5i = ?

sunisround:

(1+5i)(1-5i), for the L part of the FOIL method I got -5i * 5i = -10i^2

Vocaloid:

hint: 5*5 = ?

sunisround:

10

Vocaloid:

* means multiply what is 5 times 5?

sunisround:

alright

sunisround:

its 25

sunisround:

bruh -_- let me try it then

Vocaloid:

good, so [1 + 5i] * [1-5i] = 1^2 - (5i)^2 = 1 - 25i^2 = ?

sunisround:

26

Vocaloid:

good so let's add that to what you already wrote earlier.

Vocaloid:

(x^2-2x-3)(x^2-x[1-5i]-x[1+5i] + 26)

sunisround:

(x^2-2x-3)(x^2-x[1-5i]-x[1+5i]+26)

sunisround:

now how do i multiply all of this?

Vocaloid:

distributive property

Vocaloid:

however

Vocaloid:

(x^2-x[1-5i]-x[1+5i]+26) ^ you can simplify this to get rid of all i terms

sunisround:

how can i do that?

Vocaloid:

start by distributing inside the parentheses

sunisround:

ok so like multiply x(1-5i)?

Vocaloid:

yes (be careful with signs)

sunisround:

ok

sunisround:

do i cancel the brackets ([]) or leave them

Vocaloid:

after you distribute you should remove the appropriate brackets

sunisround:

so for the first one is it -x+5ix?

Vocaloid:

yup keep going

sunisround:

ok let me keep going

sunisround:

(x^2-x+5ix-x-5ix+26)

Vocaloid:

good, and do you see anything you can cross/cancel out?

sunisround:

ok let me try it

sunisround:

so i got (x^2-2x+26)

Vocaloid:

good, now you put that back together with what we had before

Vocaloid:

(x^2 - 2x - 3)*(x^2 - 2x + 26) then distribute

sunisround:

ok let me try it

sunisround:

x^2(x^2-2x+26)-2x(x^2-2x+26)-3(x^2-2x+26)?

Vocaloid:

yeah, but keep distributing

Vocaloid:

I think it's better if you go term by term instead of lumping (x^2-2x+26) together

sunisround:

how?

Vocaloid:



Vocaloid:

start by distributing according to these arrows

sunisround:

ok but its basically the same thing right?

Vocaloid:

yeah i guess

sunisround:

alright i like your strategy because this will not fit on my test paper

sunisround:

i am still here i am just slow for this because i am learning something new ok be patient please

Vocaloid:

no need to apologize ^^

sunisround:

so is it x^4-4x^3+37x^2-46x-78?

Vocaloid:

that 37 should be a 27 but other than that good job ^^

sunisround:

wait a 27?

sunisround:

oh yes

sunisround:

i know why 30 -3

sunisround:

ok thank you very much that is the final result right?

Vocaloid:

yup

sunisround:

how do i gratify you is there a way to give you points or something?

Vocaloid:

you can click "best response" if you'd like c;

sunisround:

where on your last comment?

Vocaloid:

it doesn't matter, any comment will work

sunisround:

ok thank you very much.